if x and y are the roots of the quadratic equation {general form}then what is the value ofx/y whole square+y/x whole square is equal to?

Grade:10

2 Answers

vikas askiitian expert
13 years ago

x,y are the roots of the quadratic eq at2 +bt +c

x+y =-b/a 7 xy =c/a

 now we have (x/y)2 + (y/x)2

               = x4 +y4 /(xy)2

               ={(x2 +y2)2 - 2(xy)2 }/(xy)2                                            {x4 +y +2(xy)2} = (x2 +y2)2

              ={ ( (x+y)2 -2xy)2 -2(xy)2 )/(xy)2                               {x2 +y2 +2xy) =(x+y)2

now substituting values of x+y and xy

 

              {(b2 -2ac)2 -2(ac)2 }/(ac)2    this is the value of (x/y)2 + (y/x)2

mycroft holmes
13 years ago